OpenStudy (anonymous):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle? (1 point) Select one: a. (x - 1)2 + (y + 2)2 = 3 b. (x + 1)2 + (y - 2)2 = 3 c. (x + 1)2 + (y + 2)2 = 9 d. (x + 1)2 + (y - 2)2 = 9

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

standard form of equation of circle : \(\large (x-h)^2 + (y-k)^2 = r^2 \) center = \(\large (h,k)\) radius = \(\large r\)

ganeshie8 (ganeshie8):

you're given, center = \(\large (h,k) = (-1, 2)\) and radius = \(\large r = 3\) simply plug them in the equation, you're done.

OpenStudy (anonymous):

ok thank you

ganeshie8 (ganeshie8):

whats ur ansur ?